### v1

# Algebraic Morse theory via Homological Perturbation Lemma with two applications Jun Chen

Abstract.
As a generalization of the classical killing-contractible-complexes lemma, we present algebraic Morse theory via homological perturbation lemma, in a form more general than existing presentations in the literature. Two-sided Anick resolutions due to E. Sköldberg are generalised to algebras given by quivers with relations and a minimality criterion is provided as well. Two applications of algebraic Morse theory are presented. It is shown that the Chinese algebra of rank is homologically smooth and of global dimension , and the minimal two-sided projective resolution of a Koszul algebra is constructed.
Discrete Morse theory has its origin in the work of K. S. Brown, R. Geoghegan [BG]. In that paper, one encountered a cell complex with one
vertex and infinitely many cells in each positive dimension. The authors, using ad hoc method,
collapsed this cell complex to a quotient complex with only two cells in each
positive dimension. K. S. Brown formalized and
applied the collapsing method scheme to groups with a rewriting system [Brown]. Motivated by differential topology, R. Forman [Forman] rediscovered and developed this theory as a discrete version of the usual smooth Morse theory, hence the name “discrete Morse theory”.
Since then, this subject has received much attention in combinatorial and computational topology; see [Forman02, Kozlov08, Knudson, Scoville] etc.
An algebraic version of discrete Morse theory has been developed by E. Sköldberg [S06], D. Kozlov [Kozlov], M. Jöllenbeck and V. Welker [JW]. It has many applications in algebra, such as
combinatorial commutative algebra [JW], cohomology of Lie algebras [S05, LV16, LV17b, LV19], Hochschild cohomology [Lopatkin, LV17a], and operad theory [DK13].
E. Sköldberg [S18] studied algebraic Morse theory from the viewpoint of the Homological Perturbation Lemma [EM, Shih, Brown64, HK, Crainic].
This paper pushes this idea further, generalising and complementing [S06, S18] as well as [Kozlov, JW] and presents an application to Chinese algebras.
The layout of this paper is as follows.
The first section contains an introduction to the Homological Perturbation Lemma (HPL) by providing some new proofs and observations which seem to be of independent interest.
In the second section we use HPL to generalise the classical killing-contractible-complexes lemma in homological algebra.
Algebraic Morse theory is presented in the third section and our definition of the key concept “Morse matching” is more general than those in
[S06, Kozlov, JW].
One of the most important advantages of algebraic Morse theory is that it provides a method to construct two-sided Anick resolutions, compared with the original one-sided version in [Anick, AG]. This is particularly useful for Hochschild cohomology.
In the fourth section, using the corresponding Gröbner-Shirshov basis theory [Green], we furnish a construction of two-sided Anick resolutions for algebras given by quivers with relations, generalising [S06, Section 3.2].
A similar resolution with a different construction of the differentials is provided by S. Chouhy and A. Solotar [CS], which might be closely related to two-sided Anick resolutions.
The fifth section contains a criterion which gives a sufficient condition when the two-sided Anick resolution is minimal; see Theorem LABEL:Thm:_minimal_criterion. We also provide a counter-example to a result that appeared in [JW].
The sixth section contains an application of algebraic Morse theory to Chinese algebras [DK92, CEKNH]. Using the
Gröbner-Shirshov basis for Chinese algebras discovered by Y. Chen and J. Qiu [CQ], we prove that a Chinese algebra of rank is homologically smooth, that is, it admits a finite length resolution by finitely generated projective bimodules, and its global dimension is ; see Theorem LABEL:Thm:_Chinese_algebras.
In last section, we consider the quadratic algebra first studied by N. Iyudu and S. Shkarin [IS] in their classification result for Hilbert series of Koszul algebras with three generators and three relations with the goal to answer two questions from the book [PP]. In [Dots], V. Dotsenko and S. R. Chowdhury constructed the one-sided Anick resolution of . We will use the algebraic Morse theory to construct the two-sided minimal resolution of through its two-sided Anick resolution.
1. Homological Perturbation Lemma
In this section, we give an introduction to the Homological Perturbation Lemma (HPL).
We first introduce some relevant notions which appear to be well known, although to our knowledge, some of them do not appear in the literature.
Definition 1.1.
A homotopy retract datum (aka. HR datum)
consists of the following
two chain maps and between complexes and , and
a homotopy between and (so ).
An HR datum is a strong quasi-isomorphism datum (aka. SQI datum) if moreover,
and are quasi-isomorphisms.
An SQI datum is a homotopy equivalence datum (aka. HE datum) if moreover,
is homotopic to .
An HE datum is a deformation retract datum (aka. DR datum), if moreover,
A DR datum is a strong deformation retract datum (aka. SDR datum), if moreover,
Remark 1.2.
DR data and SDR data exist even from the beginning of HPL [EM, Shih, Brown64].
The name HE data appeared in the work of M. Crainic [Crainic] with a different meaning, namely that which we call SQI data.
A DR datum can be always modified to an SDR datum, as explained in [LS].
We now introduce a condition which lies at the heart of the HPL.
A perturbation of Datum (1) is a graded map on of the same degree as such that is a new differential. We call it small, if is invertible. In this case, put
and consider a new perturbed datum
with
Remark 1.3.
In a heuristic manner,
The following result summarizes known versions of the HPL.
Theorem 1.4(Homological Perturbation Lemma).
Assume that is a small perturbation of Datum (1). If Datum (1) is an HR datum (resp. SQI datum, HE datum, SDR datum), then so is
Datum (2).
Remark 1.5.
It is well known that Theorem 1.4 does NOT hold for DR data; see for example, [Crainic, 2.3 Remarks (i)].
The SDR version is the first version of the HPL that appeared in the literature [EM, Shih, Brown64]; the HR version is essentially contained in the proof of the SDR version, in particular, we have
the HE version was first shown in [HK]; the SQI version is contained in [Crainic]. For a detailed historic account, see [Hue].
We will present a direct proof for the HE version compared with the original proof of [HK], which used a mapping cylinder construction to deduce the HE version from the SDR version, and we will also provide a streamlined proof for the SQI version. In the course of the proofs, we add some which appear to be new observations.
In the proofs of this section, in order to facilitate the understanding of the reader, we often underline some terms indicating that the terms will be changed in the next equality.
For the rest of this section, assume that we are given an HR datum (1) which is
endowed with a small perturbation .
The following three equalities are crucial in the proof of the HR version of HPL.
Lemma 1.6(see [Crainic, Main Perturbation Lemma]).
We have the following equalities:
From now on we assume both the HR version and the SDR version of HPL. We are going to deduce the SQI and the HE version of HPL from these assumptions.
The following interesting observations are useful in the proof of the SQI and the HE version of HPL.
Lemma 1.7.
The maps
and
are chain maps, that is,
Proof
We will show Equality (6) and Equality (7) can be proved similarly.
In fact, we have